# A Review of Graph-Powered Data Quality Applications for IoT Monitoring Sensor Networks

### v1

@@ 4.2 Edge Level
In short, although there are approaches based on optimization and GNNs, but for sensor networks, it is still very common to establish the network topology from the geodesic distances between sensors. For instance, Do et al. [136] establish the graph adjacency matrix using the geodesic distances of sensors and locations for air quality sensor networks. Nonetheless, data-driven approaches can learn tailored graphs that better represent the implicit relationship between sensor nodes as well as graphs tailored to suit a specific task or application. These data-driven graphs can improve the performance of further applications, e.g., missing value imputation [24, 137]. Convex optimization problems can provide an informative alternative to GNNs that learn the graph representation for a specific application.
@@ 4.3 Node Level
Let’s define a node-level task as finding the map $f_{\mathcal{G}}{:\;}\mathbb{R}^{N_{1}}{\rightarrow\;}\mathbb{R}^{N_{2}}$ such that $\mathbf{y}{=}f_{\mathcal{G}}(\mathbf{X})$ where $\mathbf{X}{\in}\mathbb{R}^{N_{1}\times T}$ is a graph signal ($T{=}1$) or a set of graph signals ($T{>}1$) and $\mathbf{y}$ is the target variable, which can be a categorical variable. For instance, binary vector in an outlier detection task where each graph node can be identified as anomalous, or a scalar-valued vector for the prediction of graph signals at different nodes. Value $N_{1}$ denotes the number of observed nodes in the graph and value $N_{2}$ denotes the number of target nodes in the graph.
4.3.1 Anomaly and Outlier Detection
@@
The most common approach for the outlier detection and localization task in monitoring sensor networks is the use of residual-based models. These models train a signal reconstruction model using data with normal patterns and then the reconstruction residuals are used to identify the outliers. Linear and nonlinear graph filters have been used in the detection of outliers, $|x_{i}-f_{\mathcal{G}_{i}}(\mathbf{x})|{>}TH$, for air quality and temperature sensor networks [22, 63]. Some of these graph filters allow a distributed implementation, e.g., the distributed nonlinear polynomial graph filter (NPGF) [66]. Moreover, GNNs naturally adapt to this setting allowing the implementation of graph autoencoders using attention mechanisms, convolution, and pooling layers [143, 140]. This problem can also be seen as the detection of outliers in multivariate time series, where the time series of the sensor network measurements are fed to a graph convolutional adversarial network [144]. State-of-the-art techniques include the use of GNNs and graph convolutional autoencoders. For instance, Miele et al. [145] propose a graph convolutional autoencoder to detect anomalies in wind monitoring sensor networks. Wu et al. [43] propose the use of spatiotemporal GCN and graph wavelet networks for the detection of malfunctioning fine particulate matter (PM2.5) sensors in a large-scale sensor network, providing a centralized and a localized solution. Other works make use of spatiotemporal graphs and recurrent GCN and graph coarsening to detect anomalous sensors using sophisticated anomaly scores [142]. Darvishi et al. [139] highlight the importance of outlier detection in IoT networks used in the development of digital twins, in particular they use a deep recurrent GCN to create virtual sensors and use the residuals to detect sensor failures. Han et al. [141] propose a GCN that is fed by the representation learned by a sparse autoencoder. Residuals are standardized providing explicit anomaly localization and the maximum score is used as anomaly metric, showing good results for water sensor networks. Besides, other methods focus on obtaining a list of outlying nodes given a set of measurements, e.g., Francisquini et al. [138] develop a graph filter to detect consistently outlying nodes.
@@
Reference Year Technique Use Case
Graph Filter Ferrer-Cid et al. [22] 2022 Volterra Graph Filter Outlier Detection (VGOD) Air quality sensor network
@@
Xiao et al. [63] 2020 Nonlinear Polynomial Graph Filter (NPGF) Temperature sensor network
GCN & GAT Darvishi et al. [139] 2023 Deep Recurrent GCN Water sensor network
Deng et al. [140] 2021 Graph Learning & Graph Attention Network (GAT) Water sensor network
Han et al. [141] 2022 Sparse Autoencoder & GCN Water Sensor Network
@@
All in all, residual-based models are commonly used for anomaly and outlier detection in sensor networks since they allow for the computation of a metric per sensor, allowing the localization of the faulty sensor. The most common techniques include the use of graph filters, signal smoothness, and GNNs. The main advantage of GNNs is their ability to model complex and heterogeneous phenomena, as their architecture can include classical layers for multivariate cases, temporal components or attention mechanisms. Nevertheless, graph filters provide an interpretable alternative that depending on the data can provide competitive performance, providing good distributed and generalizable approaches, section 5.1. Although not common, graph clustering techniques could also be employed to detect anomalous sensors in a sensor network given that they have been used successfully in other fields [146]. Table 3 summarizes different graph-based models for outlier detection at the node level.
@@ 4.3.2 Missing Value Imputation
In sensor networks, losses tend to occur, either due to sensor, communication, or storage failures. Therefore, it is common for measurements to present losses, i.e., that at time $t$, only a subset of sensors $\mathcal{M}{\subset}\mathcal{V}$ collect measures. Given its importance, missing value imputation has been a very active field of research, leading to the creation of different imputation models. This task is tackled at the node level since the main goal is to infer the graph signal components at different nodes, e.g., imputing the measurements of certain sensors in the network. The imputation of missing values in sensor networks is also referred to as matrix completion, graph signal reconstruction, or multivariate time series imputation or forecasting, among others.
@@
Reference Year Technique Use Case
Smoothness Betancourt et al. [147] 2023 Random Forest & Graph Signal Smoothing Air quality sensor network
Ferrer-Cid et al. [137] 2022 Graph Learning & Signal Smoothness Air quality sensor network
@@
Mondal et al. [150] 2022 Sobolev Norm Matrix Completion Weather monitoring sensor network
@@
GCN Castro et al. [152] 2023 Time GNN - GCN and Smoothness Regularization Air quality sensor network
Chen et al. [153] 2022 Adaptive GCN Air quality sensor network
Spinelli et al. [154] 2020 Adversarial Graph Denoising Autoencoder Air quality sensor network
@@
Zhang et al. [156] 2023 Spatiotemporal Variational Autoencoder Weather monitoring sensor network
Other GNN Cini et al. [157] 2022 Graph Recurrent Imputation Network (GRIN) Smart meter network
@@
Zhang et al. [159] 2023 Graph Laplacian Pyramid Network (GLPN) Solar power sensor network
Models based on matrix completion are transductive by nature, use a matrix of spatiotemporal measurements, $\mathbf{X}{\in}\mathbb{R}^{N\times T}$, with missing values, and their objective is to impute the missing entries of the matrix via the minimization of a specific matrix norm or regularization strategy. Graph-based matrix completion algorithms have been proposed [148, 160]. Sobolev signal smoothness has also been used as regularization term to complete the missing values in weather sensor networks [150]. Other frameworks jointly learn the graph and complete the missing data via matrix completion using the l2,2 and nuclear norms [24]. Besides, matrix factorization techniques have also been used for missing value imputation by using graph kernels to model spatial relationships [149]. Figure 13 shows a missing value imputation example for an air quality monitoring network, where first a graph is learned from the data and then a graph-based matrix completion method is used to impute the missing data. There also exist models that tackle the missing value imputation problem using graph signal reconstruction techniques based on the signal smoothness criterion, with applications in air quality sensor networks [137, 147]. For instance, Qiu et al. [151] optimize the smoothness of time-varying graph signals for the imputation of PM2.5 sensor measurements.
Figure 13: Example of a missing value imputation scenario in a $O_{3}$ monitoring network in Barcelona area, Spain. First, a graph Laplacian matrix $\mathbf{L}$ is learned from the data using a smoothness-based criterion [55]. Then, a matrix completion model based on Sobolev smoothness is applied to fill the data gaps, resulting in an imputation R2 of 0.91 [62].
@@
GNNs have been widely used for imputing multivariate time series, which can be interpreted as a set of graph signals collected by sensor networks. Examples of GNN architectures used include gated GNNs [161], GCNs [153, 159] or adversarially-trained GCNs [154], graph recurrent imputation network (GRIN) [157], inductive graph neural networks which provide generalizable architectures [155], GNNs with sparse spatiotemporal attention mechanisms [158], or more tailored architectures such ash the graph Laplacian pyramid network (GLPN) [159] which proposes a signal energy maintenance step to refine the imputation. Zhang et al. [156] propose a spatiotemporal variational autoencoder that makes use of GCN to perform data imputation in an IoT scenario. For instance, the most common approach for matrix completion using GNNs is to train a GNN in the form of an autoencoder so that the reconstruction error over the non-missing entries of the matrix of observations is minimized [154]. It is worth noting that ground-truth values for missing entries are not available, therefore, loss functions can be computed over observed measurements or surrogate objective functions, e.g., artificially added missings or forecasting [157]. Table 4 summarizes the different missing value imputation methods described above. Even though convex models based on GSP have been proven useful for missing value imputation, GNNs can provide superior performance for large data sets given their ability to model highly complex functions.
@@
All in all, graph topologies provide an explicit approach to modeling the spatiotemporal relationships between sensors, which are exploited by graph-based models. It is worth mentioning that the reconstruction of a sensor signal, i.e., the creation of a virtual sensor for that sensor, can be used to impute measurements in case of losses and vice versa, a scenario seen in the next section 4.3.3. Most effective techniques include matrix completion methods that pose a graph smoothness-based optimization method to complete graph signals. These techniques are known to be scalable for large sensor networks making use of advanced optimization techniques, e.g., alternating direction method of multipliers (ADMM). As mentioned in the previous section, the advantage of GNNs is that they can include a large variety of components to deal with highly complex functions, e.g., spatiotemporal components or attention mechanisms.
@@ 4.3.3 Virtual Sensing
Virtual sensing or the creation of virtual sensors (also known as soft sensors) is a widespread application in monitoring sensor networks, where estimates for a given phenomenon and location are produced without having a physical sensor, this is the major difference with the missing value imputation task. Common application cases include the creation of virtual sensors to impute missing values, replace readings from faulty sensors, or detect faulty sensors. Therefore, we are estimating the measurements in a given node $x_{vs}$ using the measurements of the other graph nodes $\mathbf{x}_{\mathcal{M}}$. This can be tackled as a graph signal reconstruction or recovery task at a node where there is no underlying physical sensor. More generally, the creation of virtual sensors can be viewed as the signal reconstruction of a node in a graph, the estimation of an unobserved parameter, or the forecasting of multivariate time series. Another example of the application of virtual sensors is the case of dynamic sensor networks where new nodes representing physical sensors and virtual sensors can join or leave the network at any time. In the machine learning field, this case is also known as out-of-sample imputation while missing value imputation corresponds to in-sample imputation [27].
@@
Reference Year Technique Use Case
Smoothness Chen et al. [61] 2015 Total Variation (TV) Air quality sensor network
Ferrer-Cid et al. [137] 2022 Graph Learning & Signal Smoothness Air quality sensor network
@@
GCN Castro et al. [152] 2023 Time GNN - GCN and Smoothness Regularization Air quality sensor network
@@
Do et al. [136] 2020 Variational GAE Matrix Completion Air quality sensor network
@@
Cini et al. [157] 2022 Graph Recurrent Imputation Network (GRIN) Smart meter sensor network
@@
Yan et al. [23] 2022 Generative Adversarial Network (GAN) & GAT Industrial IoT network
Zhang et al. [159] 2023 Graph Laplacian Pyramid Network Solar power sensor network
@@
Virtual sensors have been created using graph signal reconstruction and GNNs methods [61, 152, 27]. For instance, Ferrer-Cid et al. [137] use the signal smoothness criterion to estimate virtual sensors for O3 sensors. In fact, graph signal reconstruction models can be used to implement virtual sensors given that the measurements at non-sampled vertices (sensors) are reconstructed using the network measurements [152, 151]. Yan et al. [23] make use of a graph attention network to create virtual sensors for rocket-mounted sensors, which are later used to detect faulty sensors. Lei et al. [149] propose a Bayesian kernelized matrix factorization model to approach the missing value imputation problem as well as the virtual sensor problem, establishing a connection with Kriging where it is intended to predict in a location where there is no sensor. Graph kernels are used to model spatial relationships. Cini et al. [157] also make use of the GRIN to create virtual sensors for an air quality sensor network. Similarly, Wu et al. [155] introduce the inductive graph neural network Kriging (IGNNK) capable of generalizing to unobserved nodes, therefore, creating virtual sensors. Moreover, matrix completion using variational graph autoencoders has been carried out to infer air pollutants at a fine-grain scale, thus, creating virtual sensors for unmonitored locations [164, 136]. Following a similar approach, Calo et al. [44] introduce a spatial air signal prediction method to estimate the pollution at unobserved locations (street level) using a mean aggregation message passing algorithm in conjunction with a GNN. Indeed, this approach can be seen as obtaining virtual sensors for those locations where the signal is predicted. In terms of phenomena estimation, which can be seen as the implementation of a virtual sensor, Jia et al. [165] develop a combination of GCN with a temporal convolutional network to estimate the quality of an industrial process. In another interesting example, Niresi et al. [167] propose a spatiotemporal data fusion GAT to perform air pollution sensor calibration, which can be viewed as obtaining a virtual sensor that fuses measurements from different sensors. In the context of precision agriculture, Vyas et al. [135] use a recurrent GNN for soil moisture forecasting from other related environmental parameters. Similarly, Custodio et al. [163] evaluate the performance of the spectral temporal GNN, which combines the GDFT and classic layers such as one-dimensional convolutions, to perform soil moisture forecasting from temperature and weather parameters. Zhao et al. [166] introduce the heterogeneous temporal GNN which is able to capture the dynamics of sensors of different nature and also include the effect of exogenous variables. Darvishi et al. [139] use GCN-implemented virtual sensors to detect outliers and replace erroneous sensor readings with those of the virtual sensor if necessary, with a focus on networks feeding digital twins.
@@
This field is continuously growing, as new graph signal reconstruction techniques, matrix completion techniques, and generalization frameworks are appearing and can be used to create virtual sensors in the context of IoT monitoring sensor networks. Table 5 lists different state-of-the-art models that have been used for the estimation of virtual sensors. As it can be noted, most virtual sensing models are based on GNNs, given their ability to model highly complex spatiotemporal data. Virtual sensors are widely used, whether for outlier detection, sensor replacement or even parameter estimation at unmonitored sites. The performance of the models depends very much on the nature of the network, whether it is heterogeneous (sensors measuring different phenomena), the amount of data available or even the smoothness of the measurements. GNNs are the most widely used technique for virtual sensing, transferability and domain generalization properties can be key to their successful application in scenarios where data may be limited.

### v2

@@ 4.2 Edge Level
In short, although there are approaches based on optimization and GNNs, for sensor networks it is still very common to establish the network topology from the geodesic distances between sensors. For instance, Do et al. [129] establish the graph adjacency matrix using the geodesic distances of sensors and locations for air quality sensor networks. Nonetheless, data-driven approaches can learn tailored graphs that better represent the implicit relationship between sensor nodes as well as graphs tailored to suit a specific task or application. These data-driven graphs can improve the performance of further applications, e.g., missing value imputation [27, 146]. Convex optimization problems can provide an informative alternative to GNNs that learn the graph representation for a specific application.
@@ 4.3 Node Level
Let’s define a node-level task as finding the map $f_{\mathcal{G}}{:\;}\mathbb{R}^{N_{1}}{\rightarrow\;}\mathbb{R}^{N_{2}}$ such that $\mathbf{y}{=}f_{\mathcal{G}}(\mathbf{X})$ where $\mathbf{X}{\in}\mathbb{R}^{N_{1}\times T}$ is a graph signal ($T{=}1$) or a set of graph signals ($T{>}1$) and $\mathbf{y}$ is the target variable, which can be categorical. For instance, binary vector in an outlier detection task where each graph node can be identified as anomalous, or a scalar-valued vector for the prediction of graph signals at different nodes. Value $N_{1}$ denotes the number of observed nodes in the graph and value $N_{2}$ denotes the number of target nodes in the graph.
4.3.1 Anomaly and Outlier Detection at Node Level
@@
The most common approach for the outlier detection and localization task in monitoring sensor networks is the use of residual-based models. These models train a signal reconstruction model using data with normal patterns and then the reconstruction residuals are used to identify the outliers. Following the GSP framework, linear and nonlinear graph filters have been used in the detection of outliers, $|x_{i}-f_{\mathcal{G}_{i}}(\mathbf{x})|{>}TH$, for air quality and temperature sensor networks [25, 69]. Some of these graph filters allow a distributed implementation, e.g., the distributed nonlinear polynomial graph filter (NPGF) [72].
GNNs naturally adapt to this application, allowing the implementation of graph autoencoders using attention mechanisms, convolution, and pooling layers [147, 148]. This problem can also be seen as the detection of outliers in multivariate time series, where the time series of the sensor network measurements are fed to a graph convolutional adversarial network [149]. State-of-the-art techniques include the use of GNNs and graph convolutional autoencoders. For instance, Miele et al. [150] propose a graph convolutional autoencoder to detect anomalies in wind monitoring sensor networks. Wu et al. [47] propose the use of spatiotemporal GCN and graph wavelet networks for the detection of malfunctioning fine particulate matter (PM2.5) sensors in a large-scale sensor network, providing a centralized and a localized solution. Other works make use of spatiotemporal graphs and recurrent GCN and graph coarsening to detect anomalous sensors using sophisticated anomaly scores [151]. Darvishi et al. [152] highlight the importance of outlier detection in IoT networks used in the development of digital twins, in particular, they use a deep recurrent GCN to create virtual sensors and use the residuals to detect sensor failures. Han et al. [153] propose a GCN that is fed by the representation learned by a sparse autoencoder. Residuals are standardized providing explicit anomaly localization and the maximum score is used as anomaly metric, showing good results for water sensor networks. Besides, other methods focus on obtaining a list of outlying nodes given a set of measurements, e.g., Francisquini et al. [154] develop a graph filter to detect consistently outlying nodes. Table 3 summarizes different graph-based models for outlier detection at the node level.
@@
Reference Year Method Use Case
GSP Graph Filter Ferrer-Cid et al. [25] 2022 Volterra Graph Filter Outlier Detection (VGOD) Air quality sensor network
@@
Xiao et al. [69] 2020 NPGF Temperature sensor network
GNN GCN & GAT Darvishi et al. [152] 2023 Deep Recurrent GCN Water sensor network
Deng et al. [148] 2021 Graph Learning & GAT
Han et al. [153] 2022 Sparse Autoencoder & GCN
@@
All in all, residual-based models are commonly used for anomaly and outlier detection in sensor networks since they allow for the computation of a metric per sensor, allowing the localization of the faulty sensor. The most common techniques include the use of graph filters, signal smoothness, and GNNs. The main advantage of GNNs is their ability to model complex and heterogeneous phenomena, as their architecture can include classical layers for multivariate cases, temporal components, or attention mechanisms. Nevertheless, graph filters provide an interpretable alternative that depending on the data can provide competitive performance, providing good distributed and generalizable approaches, section 5.2.1. Although not common, graph clustering techniques could also be employed to detect anomalous sensors in a sensor network given that they have been used successfully in other fields [155].
@@ 4.3.2 Missing Value Imputation
Sensor networks are prone to losses, either due to sensor, communication or storage failures. Therefore, it is common for measurements to present losses, i.e., at time $t$, only a subset of sensors $\mathcal{M}{\subset}\mathcal{V}$ collect measures. Given its importance, missing value imputation has been a very active field of research, leading to the creation of different imputation models. This task is tackled at the node level since the main goal is to infer the graph signal components at different nodes, e.g., imputing the measurements of certain sensors in the network. The imputation of missing values in sensor networks is also referred to as matrix completion, graph signal reconstruction, or multivariate time series imputation or forecasting, among others.
Models based on matrix completion are transductive by nature, use a matrix of spatiotemporal measurements, $\mathbf{X}{\in}\mathbb{R}^{N\times T}$, with missing values, and their objective is to impute the missing entries of the matrix via the minimization of a specific matrix norm or regularization strategy. GSP-based matrix completion algorithms have been proposed where signal smoothness metrics, e.g., Sobolev smoothness, are optimized [156, 157, 158]. Other frameworks jointly learn the graph and complete the missing data via matrix completion using the $l_{2,2}$ and nuclear norms [27]. Besides, matrix factorization techniques have also been used for missing value imputation by using graph kernels to model spatial relationships [159]. There also exist models that tackle the missing value imputation problem using graph signal reconstruction techniques based on the signal smoothness criterion, with applications in air quality sensor networks [146, 160]. For instance, Qiu et al. [161] impute PM2.5 sensor measurements by optimizing the time-varying smoothness of graph signals.
@@
Reference Year Method Use Case
GSP Smoothness Betancourt et al. [160] 2023 Random Forest & Graph Signal Smoothing Air quality sensor network
Ferrer-Cid et al. [146] 2022 Graph Learning & Signal Smoothness
@@
Mondal et al. [158] 2022 Sobolev Norm Matrix Completion Weather monitoring network
@@
GNN GCN Castro et al. [162] 2023 Time GNN - GCN and Smoothness Regularization Air quality sensor network
Chen et al. [163] 2022 Adaptive GCN
Spinelli et al. [164] 2020 Adversarial Graph Denoising Autoencoder
@@
Zhang et al. [166] 2023 Spatiotemporal Variational Autoencoder Weather monitoring network
Other GNN Cini et al. [167] 2022 GRIN Smart meter network
@@
Zhang et al. [169] 2023 GLPN Solar power sensor network
@@
In terms of GNNs, these have been widely used for imputing multivariate time series, which can be interpreted as a set of graph signals collected by sensor networks. The most common approach for matrix completion using GNNs is to train a GNN in the form of an autoencoder so that the reconstruction error over the non-missing entries of the matrix of observations is minimized [164]. It is worth noting that ground-truth values for missing entries are not available, therefore, loss functions can be computed over observed measurements or surrogate objective functions, e.g., artificially introduced missings or forecasting [167]. Examples of GNN architectures used include gated GNNs [170], GCNs [163, 169] or adversarially-trained GCNs [164]. For instance, Chen et al. [163] develop a bidirectional recurrent GNN based on graph convolutional layers to perform matrix completion and online imputations in environmental data sets. Similarly, Zhang et al. [166] propose a spatiotemporal variational autoencoder composed of graph convolution layers and recurrent units to impute missing values of multiple pollutants in environmental data sets. Zhang et al. [169] introduce the graph Laplacian pyramid network (GLPN) [169] which proposes a signal’s Dirichlet energy maintenance step to refine the imputation. Other GNNs have been used for missing value imputation like graph recurrent imputation network (GRIN) [167], inductive graph neural networks that provide generalizable architectures to graphs of different topologies [165], or GNNs with sparse spatiotemporal attention mechanisms [168]. Table 4 summarizes the different missing value imputation methods described above. Even though convex models based on GSP have been proven useful for missing value imputation, GNNs can provide superior performance for large data sets given their ability to model highly complex functions.
Figure 14: Example of a missing value imputation scenario in a $O_{3}$ monitoring network in the Barcelona area, Spain. A matrix completion model based on the Sobolev smoothness and a learned graph is applied to fill the data gaps, resulting in an imputation R2 of 0.91.
Real-world use case: Figure 14 shows a missing value imputation example for an air quality monitoring network. The IoT monitoring platform measures tropospheric ozone (O3) at 8 sites in the Barcelona area, Spain. Firstly, a data-driven graph (Laplacian matrix $\mathbf{L}$) is learned from the data matrix $\mathbf{X}{\in}\mathbb{R}^{N\times T}$, where $N$ is the number of sensors and $T$ the number of available measurements, using the methodology described in [61]. Then, matrix $\mathbf{Y}$ represents the collected network measures for a given time period, where there are some data gaps or missing values. To impute the missing values, a Sobolev smoothness matrix completion approach is used [158]. The imputation results in an R2=0.91. Other graph-based missing value imputation approaches could be utilized, e.g., other signal smoothness-based criteria or GCNs making use of the learned shift matrix $\mathbf{L}$. Online variants or small data batches could be leveraged to impute missing values in real time. In similar real-world use cases, the adaptive GCN proposed by Chen et al. [163] was compared with the GRIN for the imputation of PM2.5, radiation, and humidity measurements in a network of over 500 sensors.
@@
All in all, graph topologies provide an explicit approach to modeling the spatiotemporal relationships between sensors, which are exploited by graph-based models. It is worth mentioning that the reconstruction of a sensor signal, i.e., the creation of a virtual sensor for that sensor, can be used to impute measurements in case of losses and vice versa, a scenario seen in the next section 4.3.3. Most effective techniques include matrix completion methods that pose a graph smoothness-based optimization method to complete graph signals. These techniques may face scalability challenges in large-scale sensor networks, e.g., advanced optimization techniques like the alternating direction method of multipliers (ADMM) have been used to ease scalability problems (see section 5.1). As mentioned in the previous section, the advantage of GNNs is that they can include a large variety of components to deal with highly complex functions, e.g., spatiotemporal components or attention mechanisms.
@@ 4.3.3 Virtual Sensing
Virtual sensing or the creation of virtual sensors (also known as soft sensors) is a widespread application in monitoring sensor networks, where estimates for a given phenomenon and location are produced without having a physical sensor, this is the major difference with the missing value imputation task. Common application cases include the creation of virtual sensors to impute missing values, replace readings from faulty sensors, or detect faulty sensors. Therefore, we are estimating the measurements in a given node $x_{vs}$ using the measurements of the other graph nodes $\mathbf{x}_{\mathcal{M}}$. This can be tackled as a graph signal reconstruction or recovery task at a node where there is no underlying physical sensor. More generally, the creation of virtual sensors can be viewed as the signal reconstruction of a node in a graph, the estimation of an unobserved parameter, or the forecasting of multivariate time series. Another example of the application of virtual sensors is the case of dynamic sensor networks where new nodes representing physical sensors and virtual sensors can join or leave the network at any time. In the ML field, this case is also known as out-of-sample imputation while missing value imputation corresponds to in-sample imputation [31].
@@
Reference Year Method Use Case
GSP Smoothness Chen et al. [67] 2015 TV Air quality sensor network
Ferrer-Cid et al. [146] 2022 Graph Learning & Signal Smoothness
@@
GNN GCN Castro et al. [162] 2023 Time GNN - GCN and Smoothness Regularization Air quality sensor network
@@
Do et al. [129] 2020
@@
Cini et al. [167] 2022 GRIN Smart meter sensor network
@@
Yan et al. [26] 2022 GAN & GAT Industrial IoT network
Zhang et al. [169] 2023 GLPN Solar power sensor network
Virtual sensors have been created using graph signal reconstruction and GNNs methods [67, 162, 31]. For instance, Ferrer-Cid et al. [146] use the signal smoothness criterion to estimate virtual sensors for O3 sensors. In fact, graph signal reconstruction models based on GSP can be used to implement virtual sensors given that the measurements at non-sampled vertices (sensors) are reconstructed using the network measurements [162, 161]. Lei et al. [159] propose a Bayesian kernelized matrix factorization model to approach the missing value imputation problem as well as the virtual sensor problem, establishing a connection with Kriging where it is intended to predict in a location where there is no sensor. Graph kernels are used to model spatial relationships.
@@
GNNs have been widely used for virtual sensing applications. Yan et al. [26] make use of a graph attention network to create virtual sensors for rocket-mounted sensors, which are later used to detect faulty sensors. Cini et al. [167] also make use of the GRIN to create virtual sensors for an air quality sensor network. Similarly, Wu et al. [165] introduce the inductive graph neural network Kriging (IGNNK) capable of generalizing to unobserved nodes, therefore, creating virtual sensors. Moreover, matrix completion using variational graph autoencoders has been carried out to infer air pollutants at a fine-grain scale, thus, creating virtual sensors for unmonitored locations [173, 129]. Following a similar approach, Calo et al. [48] introduce a spatial air signal prediction method to estimate the pollution at unobserved locations (street level) using a mean aggregation message passing algorithm in conjunction with a GNN. Indeed, this approach can be seen as obtaining virtual sensors for those locations where the signal is predicted. In terms of phenomena estimation, which can be seen as the implementation of a virtual sensor, Jia et al. [174] develop a combination of GCN with a temporal convolutional network to estimate the quality of an industrial process. In another interesting example, Niresi et al. [176] propose a spatiotemporal data fusion GAT to perform air pollution sensor calibration, which can be viewed as obtaining a virtual sensor that fuses measurements from different sensors. In the context of precision agriculture, Vyas et al. [145] use a recurrent GNN for soil moisture forecasting from other related environmental parameters. Similarly, Custodio et al. [172] evaluate the performance of the spectral temporal GNN, which combines the GDFT and classic layers such as one-dimensional convolutions, to perform soil moisture forecasting from temperature and weather parameters. Zhao et al. [175] introduce the heterogeneous temporal GNN, which can capture the dynamics of different sensors and include the effect of exogenous variables. Darvishi et al. [152] use GCN-implemented virtual sensors to detect outliers and replace erroneous sensor readings with those of the virtual sensor if necessary, with a focus on networks feeding digital twins.
@@
This field is continuously growing, as new graph signal reconstruction techniques, matrix completion techniques, and generalization frameworks are appearing and can be used to create virtual sensors in the context of IoT monitoring sensor networks. Table 5 lists different state-of-the-art models that have been used for the estimation of virtual sensors. As it can be noted, most virtual sensing models are based on GNNs, given their ability to accommodate highly complex spatiotemporal relationships [17]. Virtual sensors are widely used, whether for outlier detection, sensor replacement, or even parameter estimation at unmonitored sites. The performance of the models depends very much on the nature of the network, whether it is heterogeneous (sensors measuring different phenomena), the amount of data available, or even the smoothness of the measurements. GNNs are the most widely used technique for virtual sensing, transferability and domain generalization properties can be key to their successful application in scenarios where data may be limited. In short, virtual sensing is a key application that allows for uninterrupted measurements and detection of possible anomalies or deviations, which are very important features in scenarios where the data feed digital twins.